# Learning What’s Missing: Attention Dispersion and EMA Stabilization in Length Generalization

### v1

@@ 4.2 Length Generalization at the Price of Less Precision
These conditions imply the conditions of Lemma 4.3 for the matrix $\mathbf{B}+\mathbf{D}$, thus showing that we have $\mathop{\mathrm{rank}}(\mathbf{B}+\mathbf{D})\geq v-1$.
@@
Let us fix $r\in\mathbb{V}$ and let $\mathbf{a},\mathbf{c}\in\mathbb{R}^{v}$ be defined by $a_{t}=\alpha_{t,r},c_{t}=d_{r,t}$ for $t\in\mathbb{V}$. Then for all distinct $t,u\in\mathbb{V}$: we have $d_{t,u}=a_{t}+c_{u}$. Moreover, Constraints (8) and (10) show that we have $d_{t,t}-a_{t}-c_{t}<0$. Therefore, Lemma 4.3 shows that we have $\mathop{\mathrm{rank}}(\mathbf{D})\geq v-1$.
@@
Figure 1: Summary of best metrics per model. We do not display ITR values as they are below $5\cdot 10^{-4}$ in all cases.
4.3 Resolving Attention Dispersion by Dropout
@@
Dataloader
input sequence size $s$ $1+\lfloor 2^{\mathcal{U}[0,4]}\rfloor$ $[2,17]$
ambient set size $v$ $s+\lfloor 2^{\mathcal{U}[0,4]}\rfloor$ $[3,33]$
@@
embedding dim $d$ $\lfloor(v-1)\mathcal{U}[1,4]\rfloor$ $[2,128]$
key dim $d_{k}$ $\lfloor\mathcal{U}[1,d]\rfloor$ $[1,128]$
value dim $d_{v}$ $\lfloor\mathcal{U}[v-1,d]\rfloor$ $[2,128]$
@@
2. moment decay $\beta_{2}$ $10^{\mathcal{U}[-1,-8]} [11$$10^{-1},1-10^{-8}$]
@@
peak learning rate $\eta$ $10^{\mathcal{U}[-5,-1]}$ $10^{-5},10^{-1}]$
warmup steps $\lfloor 10^{\mathcal{U}[-2,6]}\rfloor$ $0,10^{6}]$
@@
Dropout
embedding dropout $\mathop{\mathrm{relu}}(\mathcal{U}[-\frac{1}{2},\frac{1}{2}])$ $0,\frac{1}{2}]$
residual dropout $\mathop{\mathrm{relu}}(\mathcal{U}[-\frac{1}{2},\frac{1}{2}])$ $0,\frac{1}{2}]$
@@
$f_{\theta}((t_{1}))=\mathbf{B}_{t_{1},:}+\mathbf{D}_{t_{1},:}\text{ and }f_{\theta}((t_{1},t_{2}))=\mathbf{B}_{t_{2},:}+\frac{\mathbf{D}_{t_{1},:}+\mathbf{D}_{t_{2},:}}{2},$
To resolve this dispersion problem, we want the model to learn larger next token logit displacements, even in the presence of weight decay. We hypothesize that increased dropout can have this effect: in training, random subnetworks learn smaller displacements that are enough to be precise on shorter sequences. Then, during inference, the smaller displacements are accumulated into larger ones, thus counteracting the dispersion effect.
4.4 Training: NLL of One-Hot Sampled Target Distribution
4.5 Mitigating Slowdown from Noisy Sampled Targets with BEMA
Figure 2: Mean dropout rates of top portions of models per generalization gap and extra dimensions.
We hypothesize that Exponential Moving Average (EMA), a general remedy for gradient noise-induced slowdown, may provide a mitigation in this setting. We use Bias-Corrected Exponential Moving Average (BEMA) [28], that we now introduce for completeness:
$\theta^{\mathrm{EMA}}_{0}$ $=\theta_{0}\text{ and }$
We call $\alpha_{n}$ a BEMA weight.
5 Random Search Experiments
5.1 Hyperparameter Distributions
We run a hyperparameter random search to see how different hyperparameter configurations influence performance, in particular length generalization. See Table 1 for the distributions we sample the hyperparameters from. We randomly sample 260 collections of dataloader hyperparameters $v,s$ and the archictectural hyperparameters $d,d_{k},d_{v}$. Then for each of these hyperparameter collection, we sample 1000 collections of hyperparameters such as learning rate or number of warmup steps that we can vectorize over. In particular, for each of the 1000 ensemble members, we sample a single collection $\rho,\kappa,\eta$ of BEMA hyperparameters.
5.2 Metrics
We introduce three metrics to measure how well a model $f_{\theta}$ solves the set complement task. Let $\mathbf{t}\in\mathbb{X}$ be an input sequence. Recall that $f_{\theta}(\mathbf{t})\in\mathbb{R}^{v}$ is the vector of predicted unnormalized next token logits to follow $\mathbf{t}$, and we let $p_{\theta}(\mathbf{t})=\mathop{\mathrm{softmax}}(f_{\theta}(\mathbf{t}))$ denote the corresponding next token probabilities.
The most important metric which we utilize to measure how closely the predicted distribution $p_{\theta}(\mathbf{t})$ approximates the uniform distribution on legal tokens $p^{*}(\mathbf{t})$ (see Equation (1)) is total variation distance (TVD):
$\mathop{\mathrm{TVD}}(\mathbf{t};\theta)=\frac{1}{2}\sum_{i=1}^{v}|p_{\theta}(\mathbf{t})_{i}-p^{*}(\mathbf{t})_{i}|.$ (16)
A more permissive metric that does not require uniformity on legal token probabilities is illegal token probability (ITP), that measures the total probability mass put on illegal tokens:
$\mathop{\mathrm{ITP}}(\mathbf{t};\theta)=\sum_{t\in[\mathbf{t}]}p_{\theta}(\mathbf{t})_{t}.$ (17)
A yet more permissive metric that only requires that the token with the highest probability—or logit—is legal is the illegal token rate (ITR):
$\mathop{\mathrm{ITR}}(\mathbf{t};\theta)=\boldsymbol{1}_{[\mathbf{t}]}(\mathop{\mathrm{argmax}}f_{\theta}(\mathbf{t}))=\begin{cases}1&\mathop{\mathrm{argmax}}(f_{\theta}(\mathbf{t}))\in[\mathbf{t}],\\ 0&\text{else}.\end{cases}$ (18)
5.3 Dataloaders, Loss and Metric Aggregation
Figure 3: Means of top quantiles of BEMA model metrics and no BEMA metrics in OthelloGPT length generalization.
Both our training and validation dataloaders output minibatches of sequences $\mathbf{T}=(\mathbf{t}_{1},\dotsc,\mathbf{t}_{N})\in\bar{\mathbb{X}}^{N}$. The difference is in the length of the sequences: in each of our training runs, we sample a training input sequence size $2\leq s<v$, and let the training dataloader output minibatches of sequences of length $s+1$. On the other hand, in each of our runs, we let the validation dataloader output minibatches of sequences of length v1v-1. In both cases, our dataloaders sample sequences uniformly.
Let $\mathbf{T}$ denote a training or validation minibatch. We follow the standard convention to aggregate the loss and the metrics not only by averaging across the minibatch entries, but also across the prefixes of the sequences. That is, the training loss on one minibatch is
$\mathop{\mathrm{NLL}}(\mathbf{T};\theta)=\frac{1}{Ns}\sum_{i=1}^{N}\sum_{s^{\prime}=1}^{s}\mathop{\mathrm{NLL}}(\mathbf{T}_{i,:s^{\prime}},t_{i,s^{\prime}};\theta)$
and if $\mu$ is one of the metrics defined in Subsection 5.2, then the corresponding validation metric reported in our experiments is
$\mu(\mathbf{T};\theta)=\frac{1}{Ns}\sum_{i=1}^{N}\sum_{s^{\prime}=1}^{s}\mu(\mathbf{T}_{i,:s^{\prime}};\theta).$
In training, we use minibatch size $N=128$, and we report validation results based on samples of $N=1024$ sequences.
5.4 Training
Following standard conventions, we initialize parameter matrices with normal distribution of std $\sigma=0.02$ and truncated at $2\sigma$. We train each ensemble of 1000 models for a maximum number of 10 000 000 AdamW [29] steps. We follow the standard practice of disabling weight decay on embedding, and norm vectors; the latter decision is ablated in [30]. We calculate validation metrics every 10 000 training steps. If there is no improvement in validation TVD of any ensemble member for 1 000 000 steps, then we stop early. For learning rate schedule, we use linear warmup, and linear decay [31]. For each of our models, we report two sets of metrics: one for the training parameters $\theta$, and one for the BEMA parameters $\theta^{\mathrm{BEMA}}$, see Subsection 4.5.
Training 260 000 models takes 1024 NVIDIA H100 NVL (96 GB) hours. We release our code upon publication, to support reproducibility, and further research.
5.5 Experiment Results
In reporting our results, we call the difference v1sv-1-s the generalization gap. Note that this quantity signifies the amount of length generalization the model has to perform: as specified in Subsection 5.3, the training input sequences have length $s$, while the validation input sequences have length v1v-1. To prove robustness of our results to hyperparameter choices, besides reporting the top metrics per model ensemble, we also report the mean metrics of the best 1% and 10%.
See Figure 1 for a summary of best metrics per model. One can see that BEMA indeed improves length generalization. Note that it does so robustly to the BEMA hyperparameters $\eta,\rho,\kappa$ (See Subsection 4.5 for definitions), as in each model we only sample a single set of BEMA hyperparameters. We do not display ITR values as they are below $5\cdot 10^{-4}$ in all cases: it is a much easier task to make a legal next token have the highest logit than to learn to output a uniform distribution on all legal next tokens.
Let us now turn to a demonstration of how dropout can boost performance across a generalization gap—if there are enough neurons so that the task can be learned with part of them shut down. Recall that (i) as per Theorem 4.2ab, the embedding dimension $d$ and value dimension $d_{v}$ both have to be at least v1v-1, and (ii) as written in Table 1, the embedding dimension $d$ is sampled from $\lfloor(v-1)\mathcal{U}[1,4]\rfloor$, and the value dimension $d_{v}$ is sampled from $\lfloor\mathcal{U}[v-1,d]\rfloor$. To display the number of extra parameters, we use the embedding dimension multiplier $\frac{d}{v-1}$, and the value dimension coefficient, that is 0 if d=v1d=v-1, and $\frac{d_{v}-v+1}{d-v+1}$ otherwise.
In Figure 2, we plot the average embedding and residual dropout rates, for the top 0.1%, 1%, and 10% of the models with and without BEMA. We see that, under substantial generalization gap, and if the embedding dimension multiplier is big enough, then better models often have increased embedding dimension. We see a less pronounced effect in the case of residual dropout.
6 Length Generalization in OthelloGPT
Figure 4: Metrics per BEMA power–EMA power pairs in OthelloGPT length generalization.
6.1 Experiment Setup
Table 2: Hyperparameter distributions for OthelloGPT
Hyperparameter Distribution
input sequence length $s$ $\lfloor\mathcal{U}[15,60)\rfloor$
attention dropout $\mathcal{U}[0,0.2]$
embedding dropout $\mathcal{U}[0,0.5]$
residual dropout $\mathcal{U}[0,0.3]$
To test how our findings generalize to a more complex setting, we perform length generalization experiments on training a GPT-1 [32] style model to output legal Othello moves via next token prediction on random Othello games. Previously, such models were studied to see if they learn a world model: the internal representations produced by such models were probed for nonlinear [6] and linear [7] representations of board state.
Besides the fixed 4 starting positions, an Othello game has $v=60$ positions. We test how well can a model trained on the first $15\leq s<60$ moves in a random game can output a uniform distribution on the next legal tokens. To measure the performance of the models, we use the same metrics as introduced in Subsection 5.3.
The original models use learned absolute positional embedding. Since that is incompatible with length generalization, we replace it by rotary positional encoding [33]. We also keep track of 10 EMA parameters $\theta^{\mathrm{EMA}}$, as defined in Subsection 4.5, with EMA power $\kappa=0.1,0.2,\dotsc,1$ and EMA lag $\rho=10$. At validation, we produce from each set of EMA parameters 10 BEMA parameters $\theta^{\mathrm{BEMA}}$ with BEMA power $\eta=0.1,0.2,\dotsc,1$. Thus, in total, we test 100 BEMA hyperparameter sets.
We sample 100 sets of input sequence length $s$ and dropout rates following Table 2. We train each model on a training set of 10 000 000 random Othello games for 1 epoch, validating every 1000 training steps. We use training minibatch size 256.
6.2 Experiment Results
Figure 5: Generalization gap to TVD without BEMA in OthelloGPT, colored by embedding dropout rate.
First of all, let’s see a summary of metrics per generalization gap in Figure 3. To indicate the robustness of the improvement BEMA brings, out of the 100 BEMA metrics, we report the best of 100, the mean of the top 10, and the mean of all the values. We can see that already the mean of the top 10 brings substantial improvement.
If we turn to metrics by BEMA hyperparameters in Figure 4, we see that the fact that in Figure 3 the mean metric over all BEMA hyperparameters was worse than the metric without BEMA was caused by the significantly worse results for EMA power values $\kappa=0.1,0.2$. In fact, BEMA performance is robust with regards hyperparameters, yielding good results for the range $0.4\leq\kappa\leq 0.9$, $0.5\leq\eta\leq 0.8$.
It turns out that in length generalization with OthelloGPT, dropout does not give a significant improvement. See for example Figure 5 for data on TVD without BEMA, and embedding dimension. It is up to further research to determine if the embedding or value dimensions were simply not big enough, or there is another effect at play. Note also the interesting detail that for generalization gap larger than 20, the gap to TVD values seem to follow a power law. This also merits futher investigation.
7 Conclusion
We introduce the Set Complement Task, that abstracts the fundamental skill of board game playing agents of detecting which positions are not yet taken. We prove that the minimal transformer models that can solve the task have to length generalize, albeit at the cost of reduced precision. Via mechanistic analysis, we uncover methods that can help mitigate said reduction. We show via random hyperparameter search that our methods are indeed effective. Finally, we show that BEMA helps length generalization in the more complex setting of OthelloGPT too.
Acknowledgments

### v2

@@ Acknowledgements
These conditions imply the conditions of Lemma A.1 for the matrix $\mathbf{B}+\mathbf{D}$, thus showing that we have $\mathop{\mathrm{rank}}(\mathbf{B}+\mathbf{D})\geq v-1$.
@@
Let us fix $r\in\mathbb{V}$ and let $\mathbf{a},\mathbf{c}\in\mathbb{R}^{v}$ be defined by $a_{t}=\alpha_{t,r},c_{t}=d_{r,t}$ for $t\in\mathbb{V}$. Then for all distinct $t,u\in\mathbb{V}$: we have $d_{t,u}=a_{t}+c_{u}$. Moreover, Constraints (9) and (11) show that we have $d_{t,t}-a_{t}-c_{t}<0$. Therefore, Lemma A.1 shows that we have $\mathop{\mathrm{rank}}(\mathbf{D})\geq v-1$.
@@
Appendix B Hyperparameter Distribution for Random Search
The following table describes the distribution from which we drew the hyperparameter configurations for our experiments.
@@ Appendix B Hyperparameter Distribution for Random Search
2. moment decay $\beta_{2}$ $10^{\mathcal{U}[-1,-5]} [11$$10^{-1},1-10^{-5}$]
@@
peak learning rate $\eta$ $10^{\mathcal{U}[-5,-2]}$ $10^{-5},10^{-2}]$
warmup steps $\lfloor 10^{\mathcal{U}[-2,4]}\rfloor$ $0,10^{4}]$
@@
Appendix C Further Experimental Results
C.1 Set Complement Task
With ambient set size $v=16,25$, we observe the same effect: including normalization of attention block output gives length generalization a big boost, and BEMA can further improve the best results.
Figure 5: Distribution of the validation TVD at the best checkpoint of each of the 1000 d=v1d=v-1, $d_{k}=1$, $d_{v}=v-1$ minimal transformers trained on the Set Complement Task with ambient set sizes $v=16,25$ and varying input sequence length $s$. In each violin plot, the left half shows the values without BEMA, and the right half with BEMA.
C.2 $\mathrm{Dyck}(k,m)$
We run further experiments on bounded Dyck languages. In all cases, we can observe similarly that peri-normalization schemes offer the best length generalization and pre-normalization the worst. We can furthermore observe a more beneficial effect of BEMA for larger numbers $k$ of types, probably due to the increased sampling noise that comes with more opening brackets being possible. Moreover, we see worse TVD with smaller maximum depth $m$, probably indicating that the EOS logits are not well calibrated. To show that, in Figure 7, we display the ROC AUC scores of the EOS tokens, showing that with peri-normalization, the models can robustly classify complete sequences.
Figure 6: Distribution of the TVD at the best checkpoints of each of the 100 2-layer, 64-dimensional, 4-headed transformers on further $\mathrm{Dyck}(k,m)$ tasks
Figure 7: Distribution of the ROC AUC of the EOS token at the best checkpoints of each of the 100 2-layer, 64-dimensional, 4-headed transformers on further $\mathrm{Dyck}(k,m)$ tasks